40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 9 00:34:34 2014, Jenne, Update, LSC, Figuring out where to do DARM->AS55 
    Reply  Tue Sep 9 14:25:46 2014, jamie, Update, LSC, Figuring out where to do DARM->AS55 
    Reply  Wed Sep 10 10:49:39 2014, Manasa, Update, LSC, Y arm green + PSL green mode overlap 
       Reply  Wed Sep 10 21:05:43 2014, Jenne, Update, LSC, Holy sensitivity, Batman! ALS_Y_outOfLoop_10Sept2014.pdfALS_XY_outOfLoop_10Sept2014.pdf
          Reply  Fri Sep 12 00:28:04 2014, ericq, Update, LSC, Holy sensitivity, Batman! ALSoutOfLoop.pdf
Message ID: 10491     Entry time: Wed Sep 10 21:05:43 2014     In reply to: 10487     Reply to this: 10497
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Holy sensitivity, Batman! 

Koji and Manasa did some work on the PSL green situation today (Koji is still writing that log post up), but I just measured the Yarm out of loop sensitivity, and WOAH. 

The beat is -11.5dBm at 42.8 MHz.  Koji said the sweet spot is around 30 MHz.  The out of loop sensitivity is 400 Hz RMS!  Something to note is that the Y beatnote still has a 20dB amplifier before going to the beatbox, but the X does not.  We had been worried about saturation issues with the X, so we took out the amplifier.  However, I might put it back if we win big like this.

Recall from elog 10462 that I had saved a reference of the out of loop noise for both X and Y, but Y was much noisier than X.  The references below are from that elog, and the new Y is in dark blue. (Edit, 9:18pm, updated plot measuring down to 0.01Hz.  This is the new reference on the ALS_outOfLoop_Ref.xml template).

 ALS_Y_outOfLoop_10Sept2014.pdf

EDIT:  (Don't worry, I'm going to measure X too, but right now the beam overlap on the camera is not good, as if something drifted after Koji and Manasa closed up the PSL table)

Touched up the alignment for X on the PSL table.  Current beatnotes are:  [Y, -13.5 dBm, 74.1 MHz], [X, -22 dBm, 13.9 MHz].  Red is the current X out of loop, and I've saved it as the new X reference on the template.

ALS_XY_outOfLoop_10Sept2014.pdf

ELOG V3.1.3-